# Preferential Attachment Trees with Vertex Death: Persistence of the Maximum Degree

### v1

@@ 4.3. Preliminary results
($\mathbb{P}\big{(}\{\mathcal{E}(u,\ell)\text{ occurs for infinitely many }\ell\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big{)}.$ (4.56)
@@
($\mathbb{P}\big{(}\{\mathcal{E}(u,\ell)\text{ occurs for infinitely many }\ell\}\big{)}\mathbb{P}\big{(}\mathcal{B}(u)<\infty,D^{(u)}=\infty\big{)}.$ (4.57)
@@
[…]
@@
[…]
@@
We then move to part $(b)$ of Theorem 4.2 and Theorem 4.4. To this end, we define for $u,v\in\mathcal{U}_{\infty}$ such that $\mathcal{B}(u)\leq\mathcal{B}(v)<\infty$, the event
@@
That is, for an individual $v$ that is born at the same time as or after $u$, we say that $v$ catches up to $u$ if at some point its offspring is at least as large as that of $u$ (not including having zero children). Heuristically, one would expect that individuals that are born ‘late’ in the branching process are unlikely to catch up to the root (if the root were to produce an infinite offspring). The following result makes this heuristic precise.
@@
Let the inter-birth times $(X(i))_{i\in\mathbb{N}}$ and the offspring random variable $D$ satisfy Assumption 4.1 and suppose that there exist $\lambda>0$ and $K\in\mathbb{N}$ such that (PHC) is satisfied. Then,
@@
$\mathbb{P}_{\mathcal{S}}{}$ $(v\text{ catches up to }\varnothing\text{ and }D^{(v)}=D^{(\varnothing)}=\infty)$ (4.62)
=($=\mathbb{P}\!\left(\{D^{(v)}=D^{(\varnothing)}=\infty\}\cap\{\mathcal{B}(v)<\infty\}\cap\{\exists j\in\mathbb{N}_{0}:\mathcal{B}(v(v_{1}+j))\leq\mathcal{B}(v_{1}+j)\}\right).$
We note that we can omit the conditioning on survival, since the event that $\varnothing$ produces an infinite offspring guarantees survival. We also need only consider children numbered $v_{1}+j$ with $j\geq 0$ in the second line, since $\varnothing$ gives birth to $v_{1}$ many children before $v$ is born. The event $\mathcal{B}(v)<\infty$ is equivalent to $D^{(v_{1}\cdots v_{\ell})}\geq v_{\ell+1}$ for all $\ell\in[m-1]$ (the event $\{D^{(\varnothing)}=\infty\}$ already implies that $v_{1}$ is born). When $D^{(v)}=\infty$, the event that $v$ gives birth to its $(v_{1}+j)^{\mathrm{th}}$ child before $\varnothing$ does is equivalent to
$\sum_{\ell=1}^{m-1}\sum_{i=1}^{v_{\ell+1}}X(v_{1}\cdots v_{\ell}i)+\sum_{i=1}^{v_{1}+j}X(vi)\leq\sum_{i=v_{1}+1}^{v_{1}+j}X(i).$ (4.63)
@@
$\mathbb{P}_{\mathcal{S}}{}$ $(v\text{ catches up to }\varnothing\text{ and }D^{(v)}=D^{(\varnothing)}=\infty)$ (4.64)
@@
We bound $\mathbb{P}\!\left(D=\infty\right)\leq\mathbb{P}\!\left(D\geq v_{1}\right)$ and apply Lemma 4.8 to arrive at the upper bound
@@
[…]
We extend the range $i\geq v_{1}+1$ in the first product to $i\geq K+1$ (as $v_{1}\geq K$) and then extend the range $v_{1}\geq K$ in the second sum to $v_{1}\geq 1$ to obtain an upper bound. As the different terms in the second product depend on distinct summation indices $v_{1},\ldots,v_{m}$, we can switch the summation and product to obtain
$\prod_{i=K+1}^{\infty}{}$ $\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]\sum_{m=1}^{\infty}\bigg{(}\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\mathrm{e}^{-\lambda\sum_{i=1}^{j}X(i)}\right]\bigg{)}^{m}$ (4.66)
[…]
where we use (4.7) in the penultimate step. Under the assumptions in the proposition, it is clear that this upper bound is finite. ∎
@@
$\mathrm{Pre}(u):=\{v\in\mathcal{U}_{\infty}:\mathcal{B}(v)\leq\mathcal{B}(u)\}\qquad\text{for }u\in\mathcal{U}_{\infty}$ (4.67)
@@
$\mathcal{P}_{m}:={}$ $\bigcup_{\begin{subarray}{c}C\subseteq\mathrm{Pre}(u)\\ |C|<m\end{subarray}}\{u\in\mathcal{U}_{\infty}:D^{(u)}=\infty\text{, }u\text{ catches up to all }v\in\mathrm{Pre}(u)\setminus C\}$ (4.68)
@@
to be the set of all individuals $u$ whose offspring diverges to infinity and at some point in time obtain a larger offspring than all but at most $m-1$ many of its predecessors. Note that this point in time need not be the same for all these predecessors. That is, it can catch up to a predecessor $v_{1}$ at time $t_{1}$ and catch up to a predecessor $v_{2}$ at time $t_{2}\neq t_{1}$. A (potential) persistent $m$-hub $u$ needs to attain the $m^{\mathrm{th}}$ largest offspring for all but a finite amount of time. As a result, there can be at most $m-1$ many predecessors of $u$ that it does not catch up to. The main ingredient for the proof of Theorems 4.2 $(b)$ and 4.4 is to ‘extend’ the result in Proposition 4.13 to the set $\mathcal{P}_{m}$, as follows.
@@
Fix $m\in\mathbb{N}$ and recall $\mathcal{P}_{m}$ from (4.68). Let the inter-birth times $(X(i))_{i\in\mathbb{N}}$ and the offspring random variable $D$ be such that Assumption 4.1 and the conditions in (PHC) are satisfied. Then, $|\mathcal{P}_{1}|$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely. If, additionally, $X(i)$ is exponentially distributed for every $i\in\mathbb{N}$, then $|\mathcal{P}_{m}|$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely for any $m\in\mathbb{N}$.
As only finitely many individuals can catch up to all but at most $m-1$ of their predecessors, and out of all pairs of these vertices one must ‘win’, as in the sense of (4.32), Theorems 4.2 $(b)$ and 4.4 then readily follow:
Proof of Theorem 4.2 $(b)$ and Theorem 4.4 subject to Proposition 4.14.
Let us first observe that combining the second condition in (PHC) with $(b)$ in Assumption 4.1 implies that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ converges almost surely. As a result, by Lemma 4.12 $(b)$,
@@
$\bigcap_{u,v\in S}\text{Win}(u,v)\cup\text{Win}(v,u)=\{\exists w\in S\ \exists T\geq 0:\deg_{t}(w)=\max_{u\in S}\deg_{t}(u)\text{ for all }t\geq T\}.$ (4.70)
As a result, with $S=\mathcal{P}_{m}$ being $\mathbb{P}_{\mathcal{S}}$-almost surely finite by Proposition 4.14, it follows that
$\mathbb{P}_{\mathcal{S}}\left(\exists w\in\mathcal{P}_{m}\ \exists T\geq 0:\deg(w,\mathrm{BP}(t))=\max_{u\in\mathcal{P}_{m}}\deg_{t}(u)\text{ for all }t\geq T\right)=1.$ (4.71)
Since a persistent $m$-hub attains the $m^{\mathrm{th}}$ largest offspring for all but a finite amount of time, it must be in $\mathcal{P}_{m}$. Hence, this proves that $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely. ∎
@@
$I(u):=\begin{cases}K&\mbox{if }D^{(u)}=\infty,\\ D^{(u)}&\mbox{if }D^{(u)}<\infty\end{cases}$ (4.72)
@@
$\mathcal{U}_{K}^{<\infty}=\mathcal{U}_{K}^{<\infty}((D^{(u)})_{u\in\mathcal{U}_{\infty}}):=\{u\in\mathcal{U}_{\infty}:u_{j}\leq I(u_{1}\cdots u_{j-1})\text{for all }j\in[|u|]\}\cup\{\varnothing\}.$ (4.73)
@@
That is, given $(D^{(u)})_{u\in\mathcal{U}_{\infty}}$, one constructs the set $\mathcal{U}_{K}^{<\infty}$ iteratively as follows. Initialise $\mathcal{U}^{(0)}:=\{\varnothing\}$ and, given $\mathcal{U}^{(i)}$ with $i\in\mathbb{N}_{0}$, we set
$\mathcal{U}^{(i+1)}:=\mathcal{U}^{(i)}\cup\bigcup_{u\in\mathcal{U}^{(i)}}\{u1,\ldots,uI(u)\}.$ (4.74)
@@
Note that, if $u\in\mathcal{U}^{(i)}$ for some $i$ and $D^{(u)}=0$, then we do not add any children $uj$ to $\mathcal{U}^{(i+1)}$. This iterative process either never terminates, in which case $\mathcal{U}^{<\infty}_{K}=\cup_{i=0}^{\infty}\mathcal{U}^{(i)}$ is infinite, or it terminates at some (random) step $M\in\mathbb{N}_{0}$, so that $\mathcal{U}^{<\infty}_{K}=\mathcal{U}^{(M)}$ is finite. We observe that, conditionally on survival of the process $(\mathrm{BP}(t))_{t\geq 0}$, the set $\mathcal{U}^{<\infty}_{K}$ contains at least one individual $u$ such that $D^{(u)}=\infty$, almost surely, since $\mathbb{P}\!\left(D=\infty\right)>0$ by Assumption 4.1.
Proof of Proposition 4.14 for $m=1$.
Recall $K$ from (PHC). We define
$T=T(K):=\inf\{t\geq 0:|\{u\in\mathcal{U}_{\infty}:D^{(u)}=\infty,\mathcal{B}(uK)\leq t\}|\geq 1\}$ (4.75)
@@
to be the stopping time at which the branching process $\mathrm{BP}$ contains at least one individual that has produced $K$ many children and that satisfies that their ‘eventual’ offspring is infinite. We let $V$ denote this individual. In case there are more than one such individual who have produced $K$ children by time $T$, then choose the first one according to their lexicographical ordering. We observe that conditionally on survival such an individual $V$ exists and that $T$ is finite, almost surely, since $\mathbb{P}\!\left(D=\infty\right)>0$. We can then write
$\mathcal{P}_{1}\subseteq\mathrm{BP}(T)\cup\{u\in\mathcal{U}_{\infty}:D^{(u)}=\infty,\ \mathcal{B}(u)>\mathcal{B}(T),\text{ $u$ catches up to all individuals in }\mathrm{Pre}(u)\}.$ (4.76)
As $T$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, it follows that $|\mathrm{BP}(T)|$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely by Lemma 4.7. It remains to show that the second set on the right-hand side has a finite size almost surely. To this end, we define
@@
$\mathcal{A}^{-}:=\mathcal{U}^{<\infty}_{K}\cap\mathrm{BP}(T)\qquad\text{and}\qquad\mathcal{A}^{+}:=\bigcup_{u\in A^{-}}\{u1,\ldots,uI(u)\}\setminus\mathcal{A}^{-},$ (4.77)

### v2

@@ 4.2. Preliminary results
$\mathbb{P}\big(\{\mathcal{E}(u,\ell)\text{ occurs for infinitely many }\ell\}\cap\{\mathcal{B}(u)<\infty,D^{(u)}=\infty\}\big).$ (4.39)
@@
$\mathbb{P}\big(\{\mathcal{E}(u,\ell)\text{ occurs for infinitely many }\ell\}\big)\mathbb{P}\big(\mathcal{B}(u)<\infty,D^{(u)}=\infty\big).$ (4.40)
@@
$\mathbb{P}_{\mathcal{S}}\left(u\text{ is a persistent $m$-hub}\right)=0,$ (4.41)
@@
$\mathbb{P}_{\mathcal{S}}\left(\exists u\in\mathcal{U}_{\infty}:u\text{ is a persistent $m$-hub}\right)=0,$ (4.42)
@@
4.4. Proving Theorem 4.1(b)
We then move to part $(b)$ of Theorem 4.1. To this end, we define for $u,v\in\mathcal{U}_{\infty}$ such that $\mathcal{B}(u)\leq\mathcal{B}(v)<\infty$, the event
@@
That is, for an individual $v$ that is born at the same time as or after $u$, we say that $v$ catches up to $u$ if at some point its offspring is at least as large as that of $u$ (not including having zero children). Heuristically, one would expect that individuals that are born ‘late’ in the branching process are unlikely to catch up to the root (if the root were to produce infinite offspring). The following result makes this heuristic precise.
@@
Recall the CTPAVD branching process $(\mathrm{BP}(t))_{t\geq 0}$ defined in (3.4), with offspring, inter-birth times, and lifetimes as defined in (the paragraph before) (3.7). Assume that Assumptions (N-E), (F-L), and (C-V) are satisfied and Assumption (F-D) is not satisfied. Then, there exists $K\in\mathbb{N}$ such that
@@
[…]
$=\mathbb{P}\!\left(\{D^{(v)}=D^{(\varnothing)}=\infty\}\cap\{\mathcal{B}(v)<\infty\}\cap\{\exists j\in\mathbb{N}_{0}:\mathcal{B}(v(v_{1}+j))\leq\mathcal{B}(v_{1}+j)\}\right).$
We note that we can omit the conditioning on survival, since the event that $\varnothing$ produces infinite offspring guarantees survival. We also need only consider children numbered $v_{1}+j$ with $j\geq 0$ in the second line, since $\varnothing$ gives birth to $v_{1}$ many children before $v$ is born. The event $\mathcal{B}(v)<\infty$ is equivalent to $D^{(v_{1}\cdots v_{\ell})}\geq v_{\ell+1}$ for all $\ell\in[m-1]$ (the event $\{D^{(\varnothing)}=\infty\}$ already implies that $v_{1}$ is born). When $D^{(v)}=\infty$, for $j\in\mathbb{N}_{0}$ the event that $v$ gives birth to its $(v_{1}+1+j)^{\mathrm{th}}$ child before $\varnothing$ does is equivalent to
$\sum_{\ell=1}^{m-1}\sum_{i=0}^{v_{\ell+1}-1}E^{(v_{1}\cdots v_{\ell})}_{i}+\sum_{i=0}^{v_{1}+j}E^{(v)}_{i}\leq\sum_{i=v_{1}}^{v_{1}+j}E^{(\varnothing)}_{i}$. (4.46)
@@
$={}$ $\mathbb{P}\!\left(\exists j\in\mathbb{N}_{0}\colon\sum_{\ell=1}^{m-1}\sum_{i=0}^{v_{\ell+1}-1}\!\!\!E^{(v_{1}\cdots v_{\ell})}_{i}+\!\sum_{i=0}^{v_{1}+j}\!E^{(v)}_{i}\leq\sum_{i=v_{1}}^{v_{1}+j}\!E^{(\varnothing)}_{i}\right)\mathbb{P}\!\left(D=\infty\right)^{2}\prod_{\ell=1}^{m-1}\mathbb{P}\!\left(D\geq v_{\ell+1}\right).$
We bound $\mathbb{P}\!\left(D=\infty\right)\leq\mathbb{P}\!\left(D\geq v_{1}\right)$ and apply Lemma 4.3 (the condition in (4.8) is satisfied for any $\lambda>0$ by choosing $K=K(\lambda)$ large, see Remark 4.6(i)(i) and $(ii)$) with
$Y=\sum_{\ell=1}^{m-1}\sum_{i=0}^{v_{\ell+1}-1}\!\!\!E^{(v_{1}\cdots v_{\ell})}_{i},\qquad X(i)=E^{(v)}_{i-1},\qquad X(i)^{\prime}=E^{(\varnothing)}_{i-1},\qquad k=v_{1}+1,$ (4.48)
to arrive at the upper bound
@@
$\leq{}$ $\sum_{m=1}^{\infty}\sum_{v_{1}=K}^{\infty}\sum_{v_{2},\ldots v_{m}=1}^{\infty}\prod_{i=v_{1}}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right].$
We observe that $\mathbb{E}\left[\exp(\lambda(E_{i}^{\prime}-E_{i}))\right]\geq 1$ by Jensen’s inequality). Hence, as $v_{1}\geq K$ we can extend the range $i\geq v_{1}$ in the first product to $i\geq K$, and also extend the range $v_{1}\geq K$ in the second sum to $v_{1}\geq 1$, to obtain an upper bound. As the different terms in the second product depend on distinct summation indices $v_{1},\ldots,v_{m}$, we can switch the summation and product to obtain
$\prod_{i=K}^{\infty}{}$ $\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\sum_{m=1}^{\infty}\bigg(\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{j-1}E_{i}\bigg)\right]\bigg)^{m}$ (4.50)
$={}$ $\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\sum_{m=1}^{\infty}\mathbb{E}\left[\sum_{j=1}^{D}\exp\bigg(-\lambda\sum_{i=1}^{j}E_{i}\bigg)\right]^{m}=\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\sum_{m=1}^{\infty}\widehat{\mu}(\lambda)^{m},$
where we use the definition of $\widehat{\mu}$, as in (2.8), in the final step. We now choose $\lambda$ large enough so that $\widehat{\mu}(\lambda)<1$, which is possible by Assumption (F-L) (as follows from (4.2)). It is then clear that the upper bound is finite, which concludes the proof. ∎
@@
$\mathrm{Pre}(u)\coloneq\{v\in\mathcal{U}_{\infty}:\mathcal{B}(v)\leq\mathcal{B}(u)\}\qquad\text{for }u\in\mathcal{U}_{\infty}$ (4.51)
@@
$\mathcal{P}_{m}\coloneq{}$ $\bigcup_{\begin{subarray}{c}C\subseteq\mathrm{Pre}(u)\\ |C|<m\end{subarray}}\{u\in\mathcal{U}_{\infty}:D^{(u)}=\infty\text{, }u\text{ catches up to all }v\in\mathrm{Pre}(u)\setminus C\}$ (4.52)
@@
to be the set of all individuals $u$ whose offspring diverges to infinity and at some point in time obtain a larger offspring than all but at most m1m-1 many of its predecessors. Note that this point in time need not be the same for all these predecessors. That is, it can catch up to a predecessor $v_{1}$ at time $t_{1}$ and catch up to a predecessor $v_{2}$ at time $t_{2}\neq t_{1}$. A (potential) persistent $m$-hub $u$ needs to attain the $m^{\mathrm{th}}$ largest offspring for all but a finite amount of time. As a result, there can be at most m1m-1 many predecessors of $u$ that it does not catch up to. The main ingredient for the proof of Theorem 4.1(b)(b) is to ‘extend’ the result in Proposition 4.8 to the set $\mathcal{P}_{m}$, as follows.
@@
Recall the CTPAVD branching process $(\mathrm{BP}(t))_{t\geq 0}$ defined in (3.4), with offspring, inter-birth times, and lifetimes as defined in (the paragraph before) (3.7). Assume that Assumptions (N-E), (F-L), and (C-V) are satisfied and Assumption (F-D) is not satisfied. Fix $m\in\mathbb{N}$ and recall $\mathcal{P}_{m}$ from (4.52). Then, $|\mathcal{P}_{m}|$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely.
As only finitely many individuals can catch up to all but at most m1m-1 of their predecessors, and out of all pairs of these vertices one must ‘win’, as in the sense of (4.18), Theorem 4.1(b)(b) readily follows:
Proof of Theorem 4.1(b)(b) subject to Proposition 4.9.
Fix $m\in\mathbb{N}$. By Lemma 4.7(b)(b),
@@
$\bigcap_{u,v\in S}\text{Win}(u,v)\cup\text{Win}(v,u)\subseteq\{\exists w\in S\ \exists T\geq 0:\deg_{t}(w)=\max_{u\in S}\deg_{t}(u)\text{ for all }t\geq T\}.$ (4.54)
As a result, with $S=\mathcal{P}_{1}$ being $\mathbb{P}_{\mathcal{S}}$-almost surely finite by Proposition 4.9, it follows that
$\mathbb{P}_{\mathcal{S}}\left(\exists w_{1}\in\mathcal{P}_{1}\ \exists T_{1}\geq 0:\deg_{t}(w_{1})=\max_{u\in\mathcal{P}_{1}}\deg_{t}(u)\text{ for all }t\geq T_{1}\right)=1.$ (4.55)
Let us call this individual the ‘winner’ of $\mathcal{P}_{1}$. Note that this individual $w_{1}$ is unique, since otherwise, if there were winners $w_{1}$ and $w_{1}^{\prime}$, then $\deg_{t}(w_{1})=\deg_{t}(w_{1}^{\prime})$ for all $t\geq\widetilde{T}$ and some time $\widetilde{T}$. But this would imply that infinitely many of their inter-birth times are of the same length, which occurs with probability zero. We can then also recursively obtain that, for any $m\geq 2$ and with $\mathcal{W}_{m}\coloneq\{w_{1},\ldots,w_{m-1}\}$,
$\mathbb{P}_{\mathcal{S}}\left(\exists w_{m}\in\mathcal{P}_{m}\setminus\mathcal{W}_{m}\ \exists T_{m}\geq 0:\deg_{t}(w_{m})=\max_{u\in\mathcal{P}_{m}\setminus\mathcal{W}_{m}}\deg_{t}(u)\text{ for all }t\geq T_{m}\right)=1.$ (4.56)
Now, for any $m\in\mathbb{N}$ and any $v\not\in\mathcal{P}_{m}$ there exist at least $m$ predecessors $v^{\prime}_{1},\ldots,v^{\prime}_{m}$ of $v$ such that $\deg_{t}(v^{\prime}_{i})>\deg_{t}(v)$ for all $t\geq\mathcal{B}(v^{\prime}_{i}1)$ and all $i\in[m]$ (that is, we do not include the time when both $v$ and $v_{i}^{\prime}$ have no children). Either $v_{i}^{\prime}\in\mathcal{P}_{m}$ for all $i\in[m]$ or there exists yet another $m$ predecessors $v^{\prime\prime}_{1},\ldots,v^{\prime\prime}_{m}$ of $v_{I}^{\prime}$ for some $I\in[m]$ such that $\deg_{t}(v_{i}^{\prime\prime})>\deg_{t}(v_{I}^{\prime})>\deg_{t}(v)$ for all $t\geq\mathcal{B}(v^{\prime\prime}_{i}1)$ and all $i\in[m]$. This procedure terminates, since the number of predecessors of $v$ is finite, so that there exist $u_{1},\ldots,u_{m}\in\mathcal{P}_{m}$ such that $\deg_{t}(u_{i})>\deg_{t}(v)$ for al $t\geq\mathcal{B}(u_{i}1)$ and for all $i\in[m]$. As a result, for any $v\not\in\mathcal{P}_{m}$ there does not exist a time $t$ such that $v$ has the $m^{\mathrm{th}}$ largest offspring among all individuals in $\mathrm{BP}(t)$.
We now show that a persistent $m$-hub exists $\mathbb{P}_{\mathcal{S}}$-almost surely for any $m\in\mathbb{N}$ by induction. Fix first $m=1$. The unique winner $w_{1}$ of $\mathcal{P}_{1}=\mathcal{P}_{1}\setminus\mathcal{W}_{1}$ (with $\mathcal{W}_{1}$ the empty set) has the largest offspring among all in $\mathcal{P}_{1}$ for all $t\geq T_{1}$ by (4.55). Since any $v\not\in\mathcal{P}_{1}$ does not attain the largest offspring for any time $t$, it follows that $w_{1}$ has the largest offspring among all individuals for all time $t\geq T_{1}$. This implies that a unique persistent $1$-hub $I^{(1),\mathrm{ct}}=w_{1}$ exists $\mathbb{P}_{\mathcal{S}}$-almost surely.
[…]
Finally, we prove (4.1). Fix two individuals $u,v\in\mathcal{U}_{\infty}$ such that $\mathcal{B}(u),\mathcal{B}(v)<\infty$ and $D^{(u)}=D^{(v)}=\infty$ and consider their number of children $\deg_{t}(u)$ and $\deg_{t}(v)$. Suppose that with positive probability there exists an increasing and diverging sequence $(t_{k})_{k\in\mathbb{N}}$ of times, such that $\deg_{t_{k}}(u)=\deg_{t_{k}}(v)$ for all $k\in\mathbb{N}$. As the number of children of $u$ and $v$ tend to infinity with $k$ almost surely conditionally on survival (and that both are born in the branching processes), this equality is equivalent to the existence of a sequence $(n_{\ell})_{\ell\in\mathbb{N}}$ such that
[…]
We can rewrite this as
$0\leq\mathcal{B}(v)-\mathcal{B}(u)+\sum_{i=0}^{n_{\ell}-1}E^{(v)}_{i}-E^{(u)}_{i}\leq E^{(u)}_{n_{\ell}}\qquad\text{for all }\ell\in\mathbb{N}.$ (4.58)
By Lemma 4.5 and Remark 4.6 the series $\sum_{i=0}^{\infty}(E^{(v)}_{i}-E^{(u)}_{i})$ converges almost surely. Hence, the distribution of the random variable $\mathcal{B}(v)-\mathcal{B}(u)+\sum_{i=0}^{\infty}(E^{(v)}_{i}-E^{(u)}_{i})$ does not contain an atom on $\mathbb{R}$ by [25, Lemma 3.2]. However, for any $\varepsilon>0$, the inequality $E^{(u)}_{n_{\ell}}<\varepsilon$ will hold for infinitely many $\ell$ almost surely by the Borel-Cantelli lemma, since
$\sum_{i=0}^{\infty}\mathbb{P}\!\left(E_{i}\geq\varepsilon\right)\leq\sum_{i=0}^{\infty}\mathrm{e}^{-(b(i)+d(i))\varepsilon}\leq\sum_{i=0}^{K}\mathrm{e}^{-(b(i)+d(i))\varepsilon}+\sum_{i=K+1}^{\infty}\frac{1}{(b(i)+d(i))^{2}}<\infty.$ (4.59)
Here, the final sum is finite and $K\in\mathbb{N}$ is such that $\mathrm{e}^{-(b(i)+d(i))\varepsilon}\leq(b(i)+d(i))^{-2}$ for all $i\geq K$, which is possible since $b(i)+d(i)$ tends to infinity, both due to Assumption (C-V).
The inequalities in (4.58) thus lead to a contradiction, so that any pair of vertices satisfies that their degrees are equal for only a finite amount of time. This directly implies (4.1), which concludes the proof. ∎
@@
$I(u)\coloneq\begin{cases}K&\mbox{if }D^{(u)}=\infty,\\ D^{(u)}&\mbox{if }D^{(u)}<\infty\end{cases}$ (4.60)
@@
$\mathcal{U}_{K}^{<\infty}=\mathcal{U}_{K}^{<\infty}((D^{(u)})_{u\in\mathcal{U}_{\infty}})\coloneq\{u\in\mathcal{U}_{\infty}:u_{j}\leq I(u_{1}\cdots u_{j-1})\text{ for all }j\in[|u|]\}\cup\{\varnothing\}.$ (4.61)
@@
That is, given $(D^{(u)})_{u\in\mathcal{U}_{\infty}}$, one constructs the set $\mathcal{U}_{K}^{<\infty}$ iteratively as follows. Initialise $\mathcal{U}^{(0)}\coloneq\{\varnothing\}$ and, given $\mathcal{U}^{(i)}$ with $i\in\mathbb{N}_{0}$, we set
$\mathcal{U}^{(i+1)}\coloneq\mathcal{U}^{(i)}\cup\bigcup_{u\in\mathcal{U}^{(i)}}\{u1,\ldots,uI(u)\}.$ (4.62)
@@
Note that, if $u\in\mathcal{U}^{(i)}$ for some $i$ and $D^{(u)}=0$, then we do not add any children $uj$ to $\mathcal{U}^{(i+1)}$. This iterative process either never terminates, in which case $\mathcal{U}^{<\infty}_{K}=\cup_{i=0}^{\infty}\mathcal{U}^{(i)}$ is infinite, or it terminates at some (random) step $M\in\mathbb{N}_{0}$, so that $\mathcal{U}^{<\infty}_{K}=\mathcal{U}^{(M)}$ is finite. We observe that, conditionally on survival of the process $(\mathrm{BP}(t))_{t\geq 0}$, the set $\mathcal{U}^{<\infty}_{K}$ contains at least one individual $u$ such that $D^{(u)}=\infty$, almost surely, when $\mathbb{P}\!\left(D=\infty\right)>0$ (which is the case when Assumption (F-D) is satisfied). We also define
$\mathcal{U}^{<\infty}\coloneq\mathcal{U}_{\infty}^{<\infty}=\{u\in\mathcal{U}_{\infty}\colon u_{j}\leq D^{(u)}\text{ for all }j\in[|u|]\}\cup\{\varnothing\}$ (4.63)
as the set of all individuals that are eventually born in $\mathrm{BP}$.
We now prove Proposition 4.8 for $m=2$. The case $m=1$ follows similarly, but requires fewer steps, whilst the difference for $m>2$ mainly lies in more notation and careful bookkeeping. As such, we decided to prove only the case $m=2$ to aid the reader. We provide a brief discussion after the proof where changes are necessary when $m=1$ or $m>2$.
Proof of Proposition 4.9 for $m=2$.
We first note that $\mathcal{P}_{2}\subseteq\mathcal{U}^{<\infty}$, since any individual that catches up to all its predecessors clearly needs to be born. We first construct a superset of $\mathcal{U}^{<\infty}$, which then informs a superset for $\mathcal{P}_{2}$. Recall $K$ from Proposition 4.8. We define
$T=T(K)\coloneq\inf\{t\geq 0:|\{u\in\mathcal{U}_{\infty}:D^{(u)}=\infty\text{ and }\mathcal{B}(uK)\leq t\}|\geq 2\}$ (4.64)
@@
to be the stopping time at which the branching process contains at least two individuals that have produced (at least) $K$ many children and whose offspring $D$ is infinite. Let $V_{1}$ and $V_{2}$ denote these two individuals, where we assume without loss of generality that $\mathcal{B}(V_{1}K)\leq\mathcal{B}(V_{2}K)=T$. Note that, conditionally on survival, such individuals exist and that $T$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, since $\mathbb{P}\!\left(D=\infty\right)>0$ by Assumption (F-D) (see [23, Lemma 5.3 and Corollary 6.4]). We then define
@@
$\mathcal{A}_{-}\coloneq\mathcal{U}_{K}^{<\infty}\cap\mathrm{BP}(T)\quad\text{and}\quad\mathcal{A}_{+}\coloneq\bigg(\bigcup_{u\in\mathcal{A}_{-}}\{u1,\ldots,uI(u)\}\bigg)\setminus\mathcal{A}_{-},$ (4.65)